OpenStudy (anonymous):

Which of the following is the solution to the equation aqrt 5 ^ 8n = 125(n + 3) ?

OpenStudy (anonymous):

sorry i wrote it wrong: Which of the following is the solution to the equation sqrt 5 ^ 8n = 125 ^ (n + 3) ?

jhonyy9 (jhonyy9):

aqrt wann to be sqrt

jhonyy9 (jhonyy9):

125 is 5 on exponent 3

OpenStudy (anonymous):

yes..do you know how to find the answer or what the answer is?

jhonyy9 (jhonyy9):

than 8n=3(n+3) 8n=3n+9 5n=9 n=9/5

OpenStudy (anonymous):

thats the right answer??

jhonyy9 (jhonyy9):

yes check it please you too

OpenStudy (anonymous):

i will! do you mind writing your work out? doesnt have to be long

jhonyy9 (jhonyy9):

so for getting one equality of two partes of equation than you need to do like in this case because we have there the number 5 on differents exponents - so we know again that 5 on exponent x will be equal with 5 on exponent y if x=y - so than if you check i have maked in base of 5 these two parts - left and right - and there i have got that 8n=3(n+3) - so after this you need to solve this equation for n and will get n= 9/5

jhonyy9 (jhonyy9):

you can prove it like exponent there is 8(9/5) = 125(9/5 +3)

jhonyy9 (jhonyy9):

these two parts need to be equales

OpenStudy (anonymous):

thanks so much!! do you know how to do this last one? The number of milligrams of a drug that remains in a patient’s system after t hours is given by the function A(t) = Iert. Don was given 300 milligrams of medicine which leaves his bloodstream at a rate of 20%. How much of the medicine remains in his system after 4 hours?

jhonyy9 (jhonyy9):

sorry good luck bye

OpenStudy (anonymous):

do u know how